### v2

@@ Appendix A Literature Review
It can be checked that the symmetric point is 00 while the dynamic range of it is $[-\tau,\tau]$. The symmetric and asymmetric components is defined by $F(W)=1$ and $G(W)=\frac{W}{\tau}$, respectively. It indicates $F_{\max}=1$. Furthermore, they assume the bounded weight saturation by assuming bounded weights, i.e., $\|W_{k}\|_{\infty}\leq W_{\max},\forall k\in[K]$ with a constant $W_{\max}<\tau$. Under this assumption, the lower bounds of response functions are given by
$\min\{H(W_{k})\}=$ $\ \min\{Q_{+}(W_{k})\odot Q_{-}(W_{k})\}=1-\left(\frac{\|W_{k}\|_{\infty}}{\tau}\right)^{2}$ (22)
$H^{\texttt{ASGD}}_{\min}=$ $\ \min\{H(W_{k})\}=1-\left(\frac{W_{\max}}{\tau}\right)^{2}.$ (23)
Convergence of Analog SGD. As we will show in Remark F.1 at the end of Appendix F, inequality (14) can be improved when the saturation never happens
@@
$\leq$ $\ \frac{4F_{\max}^{2}}{H^{\texttt{TT}}_{\min}}\sqrt{\frac{(f(W_{0})-f^{*})\sigma^{2}L}{K}}+2F_{\max}\sigma^{2}\times\frac{1}{K}\sum_{k=0}^{K-1}\left.\left\|\frac{G(W_{k})}{\sqrt{F(W_{k})}}\right\|^{2}_{\infty}\right/\min\{H(W_{k})\}$ (25)
@@
which is exactly the result in Wu et al. (2024a).
Convergence of Tiki-Taka. It is shown that a non-zero $\gamma$ in (16) improves the training accuracy Gokmen & Haensch (2020). However, Wu et al. (2024a) considers a special case with $\gamma=0$ while this paper considers a non-zero $\gamma$. As we will discuss latter in this section, different $\gamma$ leads to different convergence behaviors of Tiki-Taka.
With the linear response, if we also assume the bounded saturation of $P_{k}$ by letting $\|P_{k}\|_{\infty}\leq P_{\max}$, the minimal average response function is given by $H^{\texttt{TT}}_{\min}=1-\left(\frac{P_{\max}}{\tau}\right)^{2}$. The upper bound in Corollary 4.4 becomes
@@
As a comparison, without introducing a non-zero $\gamma$, Wu et al. (2024a) shows that convergence rate of Tiki-Taka is only
@@
Even though it is not a completely fair comparison since two paper relies on different assumptions, it is still worthy to compare the analysis in two paper. Wu et al. (2024a) assumes the noise should be non-zero, i.e. $[{\mathbb{E}}_{\xi}[|\nabla f(W;\xi)|]]_{i}\geq c_{\texttt{noise}}\sigma,\forall i\in[D]$ holds for a non-zero constant $c_{\texttt{noise}}$. Instead, this paper does not make this assumption but assumes that the objective is strongly convex. As mentioned in Section 4, the strong convexity is introduced only to ensure the existence of $P^{*}(W_{k})$. Therefore, we believe it can be relaxed, and the convergence rate can remain unchanged, which is left as a future work. Taking that into account, we believe the comparison can provide insight of how the non-zero $\gamma$ improves the convergence rate of Tiki-Taka.
Why does non-zero $\gamma$ improve the convergence rate of Tiki-Taka? As discussed in Section 4, $P_{k}$ is interpreted as a residual array that optimizes (15). In the ideal setting that $F(W)=1$ and $G(W)=0$, it can be shown that $P_{k}$ converges to $P^{*}(W_{k})$ if $W_{k}$ is fixed and $P_{k}$ is kept updated, even though the $W_{k}\neq W^{*}$ (hence $\nabla f(W_{k})\neq 0$).
Instead, without a non-zero $\gamma$, Wu et al. (2024a) points out that $P_{k}$ is an approximation of clear gradient by showing
@@
[…]
@@
Appendix C Useful Lemmas and Proofs
C.1 Lemma C.1: Properties of weighted norm
Lemma C.1.
@@ Appendix C Useful Lemmas and Proofs
C.2 Lemma C.2: Properties of weighted norm
A direct property from Definition 2.2 is that all $q_{+}(\cdot)$, $q_{-}(\cdot)$, and $F(\cdot)$ are bounded, as guaranteed by the following lemma.
Lemma C.2.
@@
C.3 Lemma C.3: Lipschitz continuity of analog update
Lemma C.3.
@@
Proof of Lemma C.3.
[…]
@@
C.4 Lemma C.4: Element-wise product error
Lemma C.4.
@@
=$\overset{(a)}{=}$ $\ C_{+}\left(\sum_{i\in[D],[U]_{i}[V]_{i}\geq 0}[U]_{i}[V]_{i}\right)-C_{-}\left(\sum_{i\in[D],[U]_{i}[V]_{i}\geq 0}|[U]_{i}[V]_{i}|\right)$
@@
Appendix D Proof of Theorem 2.1: Error from Pulse Update
See 2.1
Proof of Theorem 2.1.
@@
Appendix E Proof of Theorem 3.1: Implicit Bias of Analog Training
In this section, we provide a full version of Theorem 3.1. Before that, we formally define the accumulated asymmetric function $R_{c}(W):{\mathbb{R}}^{D}\to{\mathbb{R}}^{D}$ element-wise. Let $R(W):=\frac{G(W)}{F(W)}$ be the asymmetric ratio and $R_{c}(W)$ is defined by
@@
which satisfies
$\nabla R_{c}(W)=R(W)\quad\text{and}\quad\nabla\left\langle\Sigma,R_{c}(W)\right\rangle=\Sigma\odot R(W).$ (43)
Since we do not further assume stronger properties for response functions, like monotonicity, it is hard to provide strong claims on the shape of $R(W)$ or $R_{c}(W)$. Here we provide the expression of $R_{c}(W)$ for the linear response functions $Q_{+}(W)=1-\frac{W}{\tau}$, $Q_{-}(W)=1+\frac{W}{\tau}$. In this case, $F(W)=1$ and $G(W)=\frac{W}{\tau}$ based on definition (6); and hence $R(W)=\frac{G(W)}{F(W)}=\frac{W}{\tau}$. Accordingly, the accumulated asymmetric function is given by
@@
$=$ $\ \frac{1}{2\tau}\|W\|_{\Sigma}^{2}+\text{const.}$.
which is a weighted $\ell_{2}$ norm regularization term. It reduces to (13) in the scalar case.
If the ratio $R(W)$ is monotonic at each coordinate, $R_{c}(W)$ reaches its minimum at $W^{\diamond}$.
Theorem E.1 (Implicit Penalty, full version of Theorem 3.1).
@@
Appendix F Proof of Theorem 3.5: Convergence of Analog SGD
See 3.5
Proof of Theorem 3.5.
The $L$-smooth assumption (Assumption 3.2) implies that
@@
Bound of the second term (a). To bound term (a) in the RHS of (57), we leverage the assumption that noise has expectation 00 (Assumption 3.3)
@@
The third term in the RHS of (58) can be bounded by variance decomposition and bounded variance assumption (Assumption 3.3)
@@
Define the saturation vector $H(W_{k})\in{\mathbb{R}}^{D}$ by
$H(W_{k}):=$ $\ {F(W_{k})^{\odot 2}-G(W_{k})^{\odot 2}}={(F(W_{k})+G(W_{k}))\odot(F(W_{k})-G(W_{k}))}$ (61)
@@
$=$ $\ -\frac{\alpha}{2F_{\max}}\|\nabla f(W_{k})\|^{2}_{H(W_{k})}\leq 0.$
@@
$=$ $\ -\frac{\alpha}{2F_{\max}}\|\nabla f(W_{k})\|^{2}_{H(W_{k})}+\frac{\alpha\sigma^{2}}{2}\left\|\frac{G(W_{k})}{\sqrt{F(W_{k})}}\right\|^{2}_{\infty}$
@@
where the last inequality leverages the bounded variance of noise (Assumption 3.3) and the fact that $F(W_{k})$ is bounded by $F_{\max}$ element-wise.
@@
$\ \mathbb{E}_{\xi_{k}}[f(W_{k+1})]\leq f(W_{k})-\frac{\alpha}{2F_{\max}}\|\nabla f(W_{k})\|^{2}_{H(W_{k})}+\alpha^{2}LF_{\max}^{2}\sigma^{2}+\frac{\alpha\sigma^{2}}{2}\left\|\frac{G(W_{k})}{\sqrt{F(W_{k})}}\right\|^{2}_{\infty}$ (65)
@@
$\mathbb{E}_{\xi_{k}}[f(W_{k+1})]\leq$ $\ f(W_{k})-\frac{\alpha}{2F_{\max}}\|\nabla f(W_{k})\|^{2}_{H(W_{k})}+\alpha^{2}LF_{\max}^{2}\sigma^{2}+\alpha\sigma^{2}\left\|\frac{G(W_{k})}{\sqrt{F(W_{k})}}\right\|^{2}_{\infty}$ (68)
@@
Reorganizing (68), taking expectation over all $\xi_{K},\xi_{K-1},\cdots,\xi_{0}$, and averaging them for $k$ from 00 to $K-1$ deduce that
$E_{K}=$ $\ \frac{1}{K}\sum_{k=0}^{K}{\mathbb{E}}[\|\nabla f(W_{k})\odot F(W_{k})+|\nabla f(W_{k})|\odot G(W_{k})\|^{2}+4\|\nabla f(W_{k})\|^{2}_{H(W_{k})}]$ (69)
@@
Remark F.1 (Tighter bound without saturation).
Assuming the saturation never happens during the training, i.e. $H(W_{k})\geq H^{\texttt{TT}}_{\min}>0$ for all $k\in[K]$, we get a tighter bound in (68) by leveraging $\left\|\nabla f(W_{k})\right\|^{2}_{H(W_{k})}\geq\min\{H(W_{k})\}\left\|\nabla f(W_{k})\right\|^{2}\geq H^{\texttt{TT}}_{\min}\left\|\nabla f(W_{k})\right\|^{2}$
$\mathbb{E}_{\xi_{k}}[f(W_{k+1})]\leq$ $\ f(W_{k})-\frac{\alpha}{2F_{\max}}\|\nabla f(W_{k})\|^{2}_{H(W_{k})}+\alpha^{2}LF_{\max}^{2}\sigma^{2}+\alpha\sigma^{2}\left\|\frac{G(W_{k})}{\sqrt{F(W_{k})}}\right\|^{2}_{\infty}$ (70)
@@
$=$ $\ \frac{4F_{\max}^{2}}{H^{\texttt{TT}}_{\min}}\sqrt{\frac{(f(W_{0})-f^{*})\sigma^{2}L}{K}}+2F_{\max}\sigma^{2}\times\frac{1}{K}\sum_{k=0}^{K-1}\left.\left\|\frac{G(W_{k})}{\sqrt{F(W_{k})}}\right\|^{2}_{\infty}\right/\min\{H(W_{k})\}.$
It exactly reduces to the result for the convergence of Analog SGD in Wu et al. (2024a) on special linear repsonse functions, as discussed in Appendix B.
Appendix G Proof of Theorem 4.2: Convergence of Tiki-Taka
This section provides the convergence guarantee of the Tiki-Taka under the strongly convex assumption. See 4.2
G.1 Main proof
Proof of Theorem 4.2.

### v3

@@ Appendix B Relation with the result in [21]
It can be checked that the symmetric point is 0 while the dynamic range of it is $\tau,\tau]$. The symmetric and asymmetric components is defined by $F(W)=1$ and $G(W)=\frac{W}{\tau}$, respectively. It indicates $F_{\max}=1$. Furthermore, they assume the bounded weight saturation by assuming bounded weights, i.e., $\|W_{k}\|_{\infty}\leq W_{\max},\forall k\in[K]$ with a constant $W_{\max}<\tau$. Under this assumption, the lower bounds of response functions are given by
$\min\{R(W_{k})\}=$ $\ \min\{Q_{+}(W_{k})\odot Q_{-}(W_{k})\}=1-\left(\frac{\|W_{k}\|_{\infty}}{\tau}\right)^{2}$ (17)
$R^{\texttt{ASGD}}_{\min}=$ $\ \min\{R(W_{k})\}=1-\left(\frac{W_{\max}}{\tau}\right)^{2}.$ (18)
Challenge of analyzing the convergence of Tiki-Taka with generic response functions. For linear response functions (16), the recursion of residual array $P_{k}$ has a special structure, where the first and the biased term can be combined
$P_{k+1}=$ $\ P_{k}-\alpha\nabla f(\bar{W}_{k};\xi_{k})-\frac{\alpha}{\tau}|\nabla f(\bar{W}_{k};\xi_{k})|\odot P_{k}$ (19)
$=$ $\ \left(1-\frac{\alpha}{\tau}|\nabla f(\bar{W}_{k};\xi_{k})|\right)\odot P_{k}-\alpha\nabla f(\bar{W}_{k};\xi_{k})$
which is a weighted average of $P_{k}$ and $\nabla f(\bar{W}_{k};\xi_{k})$. Consequently, $P_{k}$ can be interpreted as an approximation of the average gradient. For this perspective, the transfer operation can be interpreted as biased gradient descent. However, given a generic $G(\cdot)$, the combination is no longer viable, bringing difficulties to the analysis.
Convergence of Analog SGD. As we will show in Remark 1 at the end of Appendix H, inequality (8) can be improved when the saturation never happens
@@
$\leq$ $\ ???$
@@
which is exactly the result in [21].
Convergence of Tiki-Taka. It is shown that a non-zero $\gamma$ in (10) improves the training accuracy [22]. However, [21] considers a special case with $\gamma=0$ while this paper considers a non-zero $\gamma$. As we will discuss latter in this section, different $\gamma$ leads to different convergence behaviors of Tiki-Taka.
With the linear response, if we also assume the bounded saturation of $P_{k}$ by letting $\|P_{k}\|_{\infty}\leq P_{\max}$, the minimal average response function is given by $R^{\texttt{RL}}_{\min}=1-\left(\frac{P_{\max}}{\tau}\right)^{2}$. The upper bound in Corollary 1 becomes
@@
As a comparison, without introducing a non-zero $\gamma$, [21] shows that convergence rate of Tiki-Taka is only
@@
Even though it is not a completely fair comparison since two paper relies on different assumptions, it is still worthy to compare the analysis in two paper. [21] assumes the noise should be non-zero, i.e. ${\mathbb{E}}_{\xi}[|\nabla f(W;\xi)|]]_{i}\geq c_{\texttt{noise}}\sigma,\forall i\in[D]$ holds for a non-zero constant $c_{\texttt{noise}}$. Instead, this paper does not make this assumption but assumes that the objective is strongly convex. As mentioned in Section 4, the strong convexity is introduced only to ensure the existence of $P^{*}(W_{k})$. Therefore, we believe it can be relaxed, and the convergence rate can remain unchanged, which is left as a future work. Taking that into account, we believe the comparison can provide insight of how the non-zero $\gamma$ improves the convergence rate of Tiki-Taka.
Why does non-zero $\gamma$ improve the convergence rate of Tiki-Taka? As discussed in Section 4, $P_{k}$ is interpreted as a residual array that optimizes $f(W_{k}+\gamma P)$. In the ideal setting that $F(W)=1$ and $G(W)=0$, it can be shown that $P_{k}$ converges to $P^{*}(W_{k})$ if $W_{k}$ is fixed and $P_{k}$ is kept updated, even though the $W_{k}\neq W^{*}$ (hence $\nabla f(W_{k})\neq 0$).
Instead, without a non-zero $\gamma$, [21] points out that $P_{k}$ is an approximation of clear gradient by showing
@@
${\leq}$ $\ \left(1-\frac{\beta}{C}\right)\|P_{k}-C\nabla f(W_{k})\|^{2}+O(\beta C^{\prime})\|\nabla f(W_{k})\|^{2}+\text{remainder}$
@@
Appendix C Dynamic of Non-ideal Analog Update
This section presents details about how to get the dynamic of analog update (3) appearing in Section˜2 and its error analysis. The primary distinction between digital and analog training is the weight update method. As discussed in Section 1, the weight update in AIMC hardware is implemented by Analog Update, which sends a series of pulses to the resistive elements.
Pulse update. Consider the response of one resistive element in one cycle, which involves only one pulse. Given the initial weight $w$, the updated weight increases or decreases by about $\Delta w_{\min}$ depending on the pulse polarity, where $\Delta w_{\min}>0$ is the response granularity determined by elements. The granularity is further scaled by a factor, which varies by the update direction due to the asymmetric update property of resistive elements. The notations $q_{+}(\cdot)$ and $q_{-}(\cdot)$ are used to denote the response functions on positive or negative sides, respectively, to describe the dominating part of the factor. In practice, the analog noise also causes a deviation of the effective factor from the response functions, referred to as cycle variation. It is represented by the magnitude $\sigma_{c}$ times a random variable $\xi_{c}$ with expectation 0 and variance $1$. Taking all of them into account, with $s\in\{+,-\}$ being the update direction, the updated weight after receiving one pulse is ${\tilde{U}}_{q}(w,s)$ where ${\tilde{U}}_{q}(\cdot,\cdot):{\mathbb{R}}\times\{+,-\}\to{\mathbb{R}}$ is the element-dependent update that implements the resistive element, which can be expressed as
${\tilde{U}}_{q}(w,s):=$ $\ w+\Delta w_{\min}\cdot(q_{s}(w)+\sigma_{c}\xi)$ (25)
The typical signal and noise ratio $\sigma_{c}/q_{s}(w)$ is roughly 5%-100% [80, 49], varied by the type of resistive elements. Furthermore, the response functions also vary by elements due to the imperfection in fabrication, called element variation (also referred to as device variation in literature [16]).
Equation (25) is a resistive element level equation. Existing work exploring the candidates of resistive elements usually reports the response curves similar to Figure˜1, [73, 52, 49]. Taking the difference between weights in two consecutive pulse cycles and adopting statistical approaches [80], all the element-dependent quantities, including $\Delta w_{\min}$, $q_{+}(\cdot)$, $q_{-}(\cdot)$ and $\sigma_{c}$, can be estimated from the response curves of the resistive elements.
Analog update implemented by pulse updates. Even though the update scheme has evolved over the years [16, 71], we discuss a simplified version, called Analog Update, to retain the essential properties. To update the weight $w$ by $\Delta w$, a series of pulses are sent, whose bit length (BL) is computed by $\operatorname{BL}:=\left\lceil\frac{|\Delta w|}{\Delta w_{\min}}\right\rceil$. After received $\operatorname{BL}$ pulses, the updated weight $w^{\prime}$ can be expressed as the function composition of (25) by $\operatorname{BL}$ times
Roughly speaking, given an ideal response $q_{+}(w)=q_{-}(w)=1$ and $\sigma_{c}=0$, $\operatorname{BL}$ pulses, with $\Delta w_{\min}$ increment for each individual pulse, incur the weight update $\Delta w$. Since the response granularity $\Delta w_{\min}$ is scaled by the response function $q_{s}(w)$, the expected increment is approximately scaled by $q_{s}(w)$ as well. Accordingly, we propose an approximate dynamic of Analog Update is given by $w^{\prime}\approx U_{q}(w,\Delta w)$, where $U_{q}(w,\Delta w)$ is defined in (3). The following theorem provides an estimation of the approximation error. It has been shown empirically that the response granularity can be made sufficiently small for updating [81, 82], implying $\Delta w_{\min}\ll\Delta w$. Therefore, we establish the error estimation of the approximation based on a small response granularity condition.
Theorem 4 (Error from discrete pulse update).
The proof of Theorem 4 is deferred to Appendix F. In Theorem 4, $|{\tilde{U}}^{\operatorname{BL}}_{q}(w,s)-U_{q}(w,\Delta w)|$ is the error between the true update and the proposed dynamic, while $|{\tilde{U}}^{\operatorname{BL}}_{q}(w,s)-w|$ is the difference between original weight and the updated one. Theorem 4 shows that the proposed dynamic dominates the update, and the approximation error is negligible when $\Delta w$ is small, which holds as $\Delta w$ always includes a small learning rate in gradient-based training.
Takeaway. Theorem 4 enables us to discuss the impact of response functions directly without dealing with element-specific details like response granularity $\Delta w_{\min}$ and cycle variation $\sigma_{c}$. Response functions are the bridge between the resistive element level equation (pulse update (25)) and algorithm level equation (dynamic of Analog Update (3)).
Appendix D Comparison of Residual Learning v2 and Tiki-Taka v2
Introducing a digital buffer, the proposed Residual Learning v2 has a similar form of Tiki-Taka v2 [23]. However, there are slight differences. Tiki-Taka v2 updates the digital buffer by
$H_{k+\frac{1}{2}}=H_{k}+\beta(P_{k+1}+\varepsilon_{k})$ (28)
which do not include a decay coefficient in front of $H_{k}$. Furthermore, Tiki-Taka v2 uses the gradient $\nabla f(W_{k};\xi_{k})$ that are solely computed on the main array $W_{k}$. Instead, Residual Learning v2 computes gradient on a mixed weight $\bar{W}_{k}=W_{k}+\gamma P_{k}$. As suggested by the ablation simulations in Section˜K.7, the training benefits from a non-zero $\gamma$.
Appendix E Useful Lemmas and Proofs
E.1 Lemma 1: Properties of weighted norm
Lemma 1.
@@ E.1 Lemma 1: Properties of weighted norm
E.2 Lemma 2: Properties of weighted norm
A direct property from Definition˜1 is that all $q_{+}(\cdot)$, $q_{-}(\cdot)$, and $F(\cdot)$ are bounded, as guaranteed by the following lemma.
Lemma 2.
@@ E.2 Lemma 2: Properties of weighted norm
E.3 Lemma 3: Lipschitz continuity of analog update
Lemma 3.
@@ E.3 Lemma 3: Lipschitz continuity of analog update
Proof of Lemma 3.
We prove for the case where $D=1$ and $S=1$, and the general case can be proven similarly. Notice that the absolute value $|\cdot|$ and vector norm $\|\cdot\|$, scalar multiplication $\times$ and element-wise multiplication $\odot$, are equivalent at that situation. We adopt both notations just for readability.
@@
E.4 Lemma 4: Element-wise product error
Lemma 4.
@@ E.4 Lemma 4: Element-wise product error
$\overset{(a)}{=}$ $\ C_{+}\left(\sum_{i\in[D],[U]_{i}[V]_{i}\geq 0}[U]_{i}[V]_{i}\right)-C_{-}\left(\sum_{i\in[D],[U]_{i}[V]_{i}\geq 0}|[U]_{i}[V]_{i}|\right)$
@@
Appendix F Proof of Theorem 4: Error from Pulse Update
See 4
Proof of Theorem 4.
@@ Appendix F Proof of Theorem 4: Error from Pulse Update
Appendix G Proof of Theorem 1: Implicit Bias of Analog Training
In this section, we provide a full version of Theorem 1. Before that, we formally define the accumulated asymmetric function $R_{c}(W):{\mathbb{R}}^{D}\to{\mathbb{R}}^{D}$ element-wise. Let $R(W):=\frac{G(W)}{F(W)}$ be the asymmetric ratio and $R_{c}(W)$ is defined by
@@ Appendix G Proof of Theorem 1: Implicit Bias of Analog Training
which satisfies $\nabla R_{c}(W)=R(W)$ and $\nabla\left\langle\Sigma,R_{c}(W)\right\rangle=\Sigma\odot R(W)$. Since we do not further assume stronger properties for response functions, like monotonicity, it is hard to provide strong claims on the shape of $R(W)$ or $R_{c}(W)$. Here we provide the expression of $R_{c}(W)$ for the linear response functions $Q_{+}(W)=1-\frac{W}{\tau}$, $Q_{-}(W)=1+\frac{W}{\tau}$. In this case, $F(W)=1$ and $G(W)=\frac{W}{\tau}$ based on definition (4); and hence $R(W)=\frac{G(W)}{F(W)}=\frac{W}{\tau}$. Accordingly, the accumulated asymmetric function is given by
@@
$=$ $\ \frac{1}{2\tau}\|W\|_{\Sigma}^{2}+\text{const.}$
which is a weighted $\ell_{2}$ norm regularization term. In the scalar case, it reduces to (44).
$\min_{W}~f_{\Sigma}(W):=f(W)+\frac{\Sigma}{2\tau}\|W\|^{2}$ (44)
It is noticeable that if the ratio $R(W)$ is monotonic at each coordinate, $R_{c}(W)$ reaches its minimum at $W^{\diamond}$. Therefore, $R_{c}(W)$ has no impact on the convergence only when the optimal point of $f(W)$ is $W^{\diamond}$.
Theorem 5 (Implicit Penalty, full version of Theorem 1).
@@
Appendix H Proof of Theorem 2: Convergence of Analog SGD
See 2
Proof of Theorem 2.
The $L$-smooth assumption (Assumption 1) implies that
@@ Appendix H Proof of Theorem 2: Convergence of Analog SGD
Bound of the second term (a). To bound term (a) in the RHS of (56), we leverage the assumption that noise has expectation 0 (Assumption 2)
@@
The third term in the RHS of (57) can be bounded by variance decomposition and bounded variance assumption (Assumption 2)
@@
Define the saturation vector $R(W_{k})\in{\mathbb{R}}^{D}$ by
$R(W_{k}):=$ $\ {F(W_{k})^{\odot 2}-G(W_{k})^{\odot 2}}={(F(W_{k})+G(W_{k}))\odot(F(W_{k})-G(W_{k}))}$ (60)
@@
$=$ $\ -\frac{\alpha}{2F_{\max}}\|\nabla f(W_{k})\|^{2}_{R(W_{k})}\leq 0.$
@@
$=$ $\ -\frac{\alpha}{2F_{\max}}\|\nabla f(W_{k})\|^{2}_{R(W_{k})}+\frac{\alpha\sigma^{2}}{2}\left\|\frac{G(W_{k})}{\sqrt{F(W_{k})}}\right\|^{2}_{\infty}$
@@
where the last inequality leverages the bounded variance of noise (Assumption 2) and the fact that $F(W_{k})$ is bounded by $F_{\max}$ element-wise.
@@
$\ \mathbb{E}_{\xi_{k}}[f(W_{k+1})]\leq f(W_{k})-\frac{\alpha}{2F_{\max}}\|\nabla f(W_{k})\|^{2}_{R(W_{k})}+\alpha^{2}LF_{\max}^{2}\sigma^{2}+\frac{\alpha\sigma^{2}}{2}\left\|\frac{G(W_{k})}{\sqrt{F(W_{k})}}\right\|^{2}_{\infty}$ (64)
@@
$\mathbb{E}_{\xi_{k}}[f(W_{k+1})]\leq$ $\ f(W_{k})-\frac{\alpha}{2F_{\max}}\|\nabla f(W_{k})\|^{2}_{R(W_{k})}+\alpha^{2}LF_{\max}^{2}\sigma^{2}+\alpha\sigma^{2}\left\|\frac{G(W_{k})}{\sqrt{F(W_{k})}}\right\|^{2}_{\infty}$ (67)
@@
Reorganizing (67), taking expectation over all $\xi_{K},\xi_{K-1},\cdots,\xi_{0}$, and averaging them for $k$ from 0 to K1K-1 deduce that
$E_{K}=$ $\ \frac{1}{K}\sum_{k=0}^{K}{\mathbb{E}}[\|\nabla f(W_{k})\odot F(W_{k})+|\nabla f(W_{k})|\odot G(W_{k})\|^{2}+4\|\nabla f(W_{k})\|^{2}_{R(W_{k})}]$ (68)
@@
Remark 1 (Tighter bound without saturation).
Assuming the saturation never happens during the training, i.e. $R(W_{k})\geq R^{\texttt{RL}}_{\min}>0$ for all $k\in[K]$, we get a tighter bound in (67) by leveraging $\left\|\nabla f(W_{k})\right\|^{2}_{R(W_{k})}\geq\min\{R(W_{k})\}\left\|\nabla f(W_{k})\right\|^{2}\geq R^{\texttt{RL}}_{\min}\left\|\nabla f(W_{k})\right\|^{2}$
$\mathbb{E}_{\xi_{k}}[f(W_{k+1})]\leq$ $\ f(W_{k})-\frac{\alpha}{2F_{\max}}\|\nabla f(W_{k})\|^{2}_{R(W_{k})}+\alpha^{2}LF_{\max}^{2}\sigma^{2}+\alpha\sigma^{2}\left\|\frac{G(W_{k})}{\sqrt{F(W_{k})}}\right\|^{2}_{\infty}$ (69)
@@
$=$ $\ \frac{4F_{\max}^{2}}{R^{\texttt{RL}}_{\min}}\sqrt{\frac{(f(W_{0})-f^{*})\sigma^{2}L}{K}}+2F_{\max}\sigma^{2}\times\frac{1}{K}\sum_{k=0}^{K-1}\left.\left\|\frac{G(W_{k})}{\sqrt{F(W_{k})}}\right\|^{2}_{\infty}\right/\min\{R(W_{k})\}.$
It exactly reduces to the result for the convergence of Analog SGD in [21] on special linear repsonse functions, as discussed in Appendix B.
Appendix I Proof of Theorem 3: Convergence of Residual Learning
This section provides the convergence guarantee of the Tiki-Taka under the strongly convex assumption. See 3
I.1 Main proof
Proof of Theorem 3.